OpenStudy (anonymous):

a and b are the lengths of the legs of a right triangle and c is the length of the hypotenuse. find b if a = 21 and c = 29. a. b = 40 b. b = 20 c. b=10 d. b=400

OpenStudy (smarty):

B, 20 21^2=441 29^2=841 A^2+441=881 A^2=400 A=20